# Anisotropy can make a moving active fluid membrane rough or crumpled

### v1

@@ C.6 RG recursion relations
$\lambda_{ij}^{{}^{\prime}}=\lambda_{ij}^{<}b^{-d-2}\xi.$
@@
$\lambda^{{}^{\prime}}=\lambda^{<}b^{\chi+z-2},$
$\overline{\lambda}^{{}^{\prime}}=\overline{\lambda}^{<}b^{\chi+z-2}.$
@@
Here, $\chi$ is the roughness exponent of an anisotropic KPZ surface. Since there is no correction to $\lambda$, $\overline{\lambda}$ so $\lambda^{<}=\lambda$ and $\overline{\lambda}^{<}=\overline{\lambda}$. We further get
$D^{<}$ $=$ $D\biggr{[}1+\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{[}\Bigl{(}\frac{2\nu^{2}+\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}+\frac{\overline{\lambda}^{2}}{\lambda^{2}}\Bigl{(}\frac{\nu^{2}+2\overline{\nu}^{2}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}-\frac{6\overline{\lambda}}{\lambda}\Bigl{(}\frac{\nu\overline{\nu}}{\nu^{2}-\overline{\nu}^{2}}\Bigl{)}\Bigr{]}\ln b\biggr{]},$ (48)
$\nu^{<}$ $=$ $\nu\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}\frac{\overline{\lambda}\overline{\nu}}{\lambda\nu}-(\frac{\overline{\lambda}}{\lambda})^{2}\Bigr{)}\ln b\biggr{]},$ (49)
$\overline{\nu}^{<}$ $=$ $\overline{\nu}\biggr{[}1-\frac{2\lambda^{2}D}{16(2\pi)(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr{(}1-\frac{\overline{\lambda}\nu}{\lambda\overline{\nu}}\Bigr{)}\ln b\biggr{]}.$ (50)
@@
Now set $b=e^{\delta l}\approx 1+\delta l$ for small $\delta l$, and define a dimensionless coupling constant by $g=\frac{\lambda^{2}D}{8\pi(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}$ and dimensionless ratios $\alpha=\frac{\overline{\nu}}{\nu}$ and $\beta=\frac{\overline{\lambda}}{\lambda}$ (see main text). These allow us to obtain the differential RG recursion relations reported in the main text. Using the flow equations of $g$ and $\alpha$ it turns out that $g$ is stable if $\beta^{2}>1$ and in this range of $\beta$, $\alpha^{*}=1/\beta$ is locus of stable fixed points (stable fixed line). Using $d=2$ and the linear theory results $z=2$, $\chi=0$ we find that $\nu_{l}\sim\nu_{0}l^{A}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}l^{B}$ and $D_{l}\sim D_{0}l^{C}$. Where,
$A[\beta,\alpha^{*}]=\frac{\beta^{2}-\alpha^{*}\beta}{2s},\;B[\beta,\alpha^{*}]=\frac{(\beta/\alpha^{*})-1}{2s},\;C[\beta,\alpha^{*}]=\frac{1}{2s}\biggr{[}\frac{(2+\alpha^{*2})+\beta^{2}(1+2\alpha^{*2})-6\beta\alpha^{*}}{1-\alpha^{*2}}\biggr{]}.$
Here, $s=\beta^{2}-1$. It turns out that at $\alpha=\alpha^{*}=\frac{1}{\beta}$, $A=B=C=\frac{1}{2s}(\beta^{2}-1)=\frac{1}{2}$ implying $\nu_{l}\sim\nu_{0}\sqrt{l}$, $\overline{\nu}_{l}\sim\overline{\nu}_{0}\sqrt{l}$ and $D_{l}\sim D_{0}\sqrt{l}$.
Appendix D Direct numerical simulations
@@
Following Ref. Moser et al. (1991), we discretize Eq. (11) by standard forward-backward differences on a square grid with lattice constant $a_{0}$. Employing the Euler algorithm with time increments $\Delta t$, we integrate Eq. (11). Grid points are labeled by ${\bf n}$, and the basis vector is denoted by ${\bf e}_{i}$, where $i=1,2$, giving the discretized form of Eq. (11)
$h_{\bf n}(t+\Delta t)$ $=h_{\bf n}(t)+\frac{\Delta t}{(a_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\nu\Bigl{[}h_{\bf n+e_{i}}(t)-2h_{\bf n}(t)+h_{\bf n-e_{i}}(t)\Bigl{]}-\frac{\lambda}{8}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}^{2}\biggl{\}}+\frac{2\overline{\nu}\Delta t}{4(a_{0})^{2}}\Bigl{[}h_{\bf n+e_{1}+e_{2}}(t)$
$-h_{\bf n+e_{1}-e_{2}}(t)+h_{\bf n-e_{1}-e_{2}}(t)-h_{\bf n-e_{1}+e_{2}}(t)\Bigl{]}-\frac{\overline{\lambda}\Delta t}{4(a_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl{]}+\sigma\sqrt{12\Delta t}R_{n}(t).$ (51)
@@
Here, $\sigma^{2}=\frac{2D}{(a_{0})^{2}}$. The variable $R$ corresponds to uniformly distributed random numbers ranging from $-\frac{1}{2}$ to $\frac{1}{2}$. The prefactor $\sigma\sqrt{12\Delta t}$ ensures that the noise possesses an equivalent second moment as Gaussian noise integrated over a time interval $\Delta t$. Equation (51) can be conveniently expressed using dimensionless quantities $\tilde{h}=\frac{h}{h_{0}}$, $\tilde{t}=\frac{t}{t_{0}}$, and $\tilde{a}_{0}=\frac{a_{0}}{x_{0}}$, where the natural units are defined as $h_{0}=\frac{\nu}{\lambda}$, $t_{0}=\frac{\nu^{2}}{\sigma^{2}\lambda^{2}}$ and $x_{0}=\sqrt{\frac{\nu^{3}}{\sigma^{2}\lambda^{2}}}$. In these units Eq. (51) becomes
$\tilde{h}_{\bf n}(\tilde{t}+\Delta\tilde{t})$ $=\tilde{h}_{\bf n}(\tilde{t})+\frac{\Delta\tilde{t}}{(\tilde{a}_{0})^{2}}\sum_{i=1}^{2}\biggl{\{}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-2\tilde{h}_{\bf n}(\tilde{t})+\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}-\frac{1}{8}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}^{2}\biggl{\}}+\frac{\alpha\Delta\tilde{t}}{2(\tilde{a}_{0})^{2}}\Bigl{[}\tilde{h}_{\bf n+e_{1}+e_{2}}(\tilde{t})$
$-\tilde{h}_{\bf n+e_{1}-e_{2}}(\tilde{t})+\tilde{h}_{\bf n-e_{1}-e_{2}}(\tilde{t})-\tilde{h}_{\bf n-e_{1}+e_{2}}(\tilde{t})\Bigl{]}-\frac{\beta\Delta\tilde{t}}{4(\tilde{a}_{0})^{2}}\prod_{i=1}^{2}\Bigl{[}\tilde{h}_{\bf n+e_{i}}(\tilde{t})-\tilde{h}_{\bf n-e_{i}}(\tilde{t})\Bigl{]}+\sqrt{12\Delta\tilde{t}}R_{n}(\tilde{t}).$ (52)
Figure 10: The variation of width ${\mathcal{W}}$ with time $t$ is presented on a log-log scale for different system sizes with two parameter sets: (a) $\alpha=0.50$, $\beta=2$, corresponding to the logarithmic rough region from RG analysis, with system sizes $L=100$, $300$, $500$, $800$, $1000$, and $1600$; (b) $\alpha=0.50$, $\beta=0.042$, corresponding to the algebraic rough region from our RG analysis, with system sizes $L=100$, $200$, $300$, $400$, and $500$. The red dashed line indicates a linear fit in the linear region on a log-log scale, yielding a slope of $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$ (see text).
For our simulations we chose $a_{0}=1$ and $\sigma=0.10$. First we discuss the numerical results for the logarithmic rough phase. For that, we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$, which corresponds to the logarithmic rough phase as predicted by our RG calculations (see main text). For a system of size $L$, we reached the saturation region (after the time after which width $\cal W$ saturates) by simulating Eq. (52) up to time $T\approx L^{2}$ in time step of $\Delta t=0.05$ as shown in Fig. 10(a) for different system sizes. Width $\cal W$ at a time $t$ was calculated using ${\cal W}=\sqrt{\langle\tilde{h}({\bf x},t)^{2}\rangle-\langle\tilde{h}({\bf x},t)\rangle^{2}}$. Figure 10(a) shows the variation of $\cal W$ with time $t$ in the growth and saturation region on a log-log scale and correspond to a single noise realization. The plot clearly shows that growth nearly saturates before reaching time $T\approx L^{2}$. To further investigate the saturation region, we extended the simulation for an additional $10^{6}$ time steps. During these $10^{6}$ time steps, we performed $N$ noise averages, with different choice of $N$ for different system sizes. For each of these time steps width ${\cal W}$ was calculated and averaged over $N$. The final saturated width ${\mathcal{W}_{\text{sat}}}$ was obtained by performing an average over $100000$ realizations of ${\cal W}_{\text{sat}}$ obtained after saturation for different system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$. For $L=100$, $300$ noise average $N$ was set to $100$. For $L=500$ $N$ was set to $50$. And for $L=800$, $1000$ and $1600$ $N$ was set to $20$. Variation of ${\cal W}_{\text{sat}}$ with $\ln L$ in the saturation region is shown in the main text Fig. 3(a).
@@
Next we discuss the numerical results for the algebraic rough phase. In this region, width ${\cal W}$ is expected to scale as $\mathcal{W}(L,t)\sim L^{\chi}f(\frac{t}{L^{z}})$ where, $f(x)$ is a dimensionless scaling function. With parameters $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, (which should correspond to the algebraic rough phase from RG analysis), we run the simulation for $L=100$, $200$, $300$, $400$ and $500$. For both of the growth and saturation region ${\cal W}$ was calculated at each time step $\Delta t=0.01$ with $N$ noise averages up to time $T$. For system size $L=100$, $200$, $300$, $400$ and $500$ the Eq. (52) was simulated up to time $T=5000$, $10000$, $15000$, $25000$ and $30000$ with time step $\Delta t=0.01$. $L=100$ to $400$ was noise averaged with $N=100$. For $L=500$, $N=50$ was taken. Refer to Fig. 10(b) for the variation of ${\cal W}$ over time on a log-log scale. The red dashed line indicates the slope of the growth region obtained through linear fitting, yielding the growth exponent $\tilde{\beta}=0.234\pm 1.644\times 10^{-5}$. Next we switch over to the saturation region of ${\cal W}$. With $T$ defined earlier for different $L$, we reached the saturation region of ${\cal W}$. See Fig. 10(b). Then we numerically solved the equation of motion (11) for further $100000$ time steps with $\Delta t=0.01$. And for each of these $100000$ time steps we have taken $N=500$ noise averages for all $L=100$ to $500$. For each of these time steps ${\cal W}$ was calculated and averaged over $N$. The final ${\mathcal{W}_{\text{sat}}}$ was calculated by averaging over 100000 realizations in the saturated regime. Variation of ${\cal W}_{\text{sat}}$ with $L$ in the saturation region is shown in Fig. 3(b) of the main text.

### v2

@@ III.6 RG recursion relations
$\lambda^{{}^{\prime}}=\lambda^{<}b^{-d-2}\xi.$
@@
$\lambda^{{}^{\prime}}=\lambda^{<}b^{\chi+z-2}.$
@@
Here, $\chi$ is the roughness exponent of an anisotropic KPZ surface. Since there is no correction to $\lambda$, so $\lambda^{<}=\lambda$. We further get
$D^{<}$ $=$ $D\biggr[1+\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{3\beta^{2}}{4}+\frac{1}{2}(\beta+2\alpha^{2})+\frac{3}{4}\Bigr)\frac{\ln b}{(\beta-\alpha^{2})}\biggr],$ (45)
$\nu_{x}^{<}$ $=$ $\nu_{x}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{1}{2}-\frac{\beta}{2}\Bigr)\ln b\biggr],$ (46)
$\nu_{y}^{<}$ $=$ $\nu_{y}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\Bigr(\frac{1}{2}-\frac{1}{2\beta}\Bigr)\ln b\biggr],$ (47)
$\overline{\nu}^{<}$ $=$ $\overline{\nu}\biggr[1-\frac{2\lambda^{2}D}{16(2\pi)(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}\ln b\biggr].$ (48)
@@
The extra factor of $2$ is multiplied in numerator of the above renormalized parameters to include contribution from negative values of $q_{y}$ i.e. ($\Lambda\leq q_{y}\leq-\Lambda/b$) in the loop integrals. See the lower strip in Fig. 5. Now set $b=e^{\delta l}\approx 1+\delta l$ for small $\delta l$, and define a dimensionless coupling constant by $g=\frac{\lambda^{2}D}{16\pi(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}}$ and dimensionless ratios $\alpha=\frac{\overline{\nu}}{\nu_{x}}$ and $\beta=\frac{\nu_{y}}{\nu_{x}}$ (see main text). These allow us to obtain the differential RG recursion relations reported in the main text.
IV Direct numerical simulations
@@
Following Ref. Moser et al. (1991), we discretize Eq. (1) by standard forward-backward differences on a square grid with lattice constant $a_{0}$. Employing the Euler algorithm with time increments $\Delta t$, we integrate Eq. (1). The grid points are labeled by $\bf n$}, and the basis vector is denoted by ${\bf e}_{i}$, where $i=1,2$, giving the discretized form of Eq. (1)
$h_{\bf n}(t+\Delta t)$ $=h_{\bf n}(t)+\frac{\Delta t}{(a_{0})^{2}}\sum_{i=1}^{2}\biggl\{\nu_{i}\Bigl[h_{\bf n+e_{i}}(t)-2h_{\bf n}(t)+h_{\bf n-e_{i}}(t)\Bigl]-\frac{\lambda}{8}\Bigl[h_{\bf n+e_{i}}(t)-h_{\bf n-e_{i}}(t)\Bigl]^{2}\biggl\}+\frac{\overline{\nu}\Delta t}{2(a_{0})^{2}}\Bigl[h_{\bf n+e_{1}+e_{2}}(t)$
$-h_{\bf n+e_{1}-e_{2}}(t)+h_{\bf n-e_{1}-e_{2}}(t)-h_{\bf n-e_{1}+e_{2}}(t)\Bigl]+\sqrt{\frac{24D\Delta t}{(a_{0})^{2}}}R_{\bf n}(t).$ (49)
@@
Here $\nu_{1}(\nu_{2})\equiv\nu_{x}(\nu_{y})$. The variable $R$ corresponds to uniformly distributed random numbers ranging from $\frac{1}{2}$ to $\frac{1}{2}$. The prefactor $\sqrt{\frac{24D\Delta t}{(a_{0})^{2}}}$ ensures that the noise possesses an equivalent second moment as Gaussian noise integrated over a time interval $\Delta t$.
@@
For our simulations we chose $a_{0}=1$. In the the speculated algebraic rough phase region, width $\cal W$} is expected to scale as $\mathcal{W}(L,t)\sim L^{\chi}f(\frac{t}{L^{z}})$ where, $f(x)$ is a dimensionless scaling function. With different set of parameters we run the simulation for $L=100$, $200$, $300$, $400$ and $500$. For both of the growth and saturation region $\cal W$} was calculated at different times with $N$ noise averages up to time $T$. For system size $L=100$, $200$, $300$, $400$ and $500$ the Eq. (49) was simulated up to time $T=5000$, $10000$, $15000$, $25000$ and $30000$ with time step $\Delta t=0.01$. $L=100$, $200$ was noise averaged with $N=500$. For $L=300$, $400$ and $500$ $N=200$, $100$ and $50$ was taken. Refer to Fig. 3 in the main text for the variation of $\cal W$} over time on a log-log scale. With $T$ defined earlier for different $L$, we reached the saturation region of $\cal W$}. Then we numerically solved the equation of motion (1) for further $100000$ time steps with $\Delta t=0.01$. And for each of these $100000$ time steps we have taken $N=500$ noise averages for all $L=100$ to $500$. For each of these time steps $\cal W$} was calculated and averaged over $N$. The final $\mathcal{W}_{\text{sat}}$} was calculated by averaging over 100000 realizations in the saturated regime. Variation of ${\cal W}_{\text{sat}}$ with $L$ in the saturation region is shown in Fig. 3(d) of the main text.